# GPT-4V(ision) is a Generalist Web Agent, if Grounded

### v1

@@ 4.2 Online Evaluation Results
Across all three settings, SeeActChoice outperforms both GPT-4 and FLAN-T5-XL. Using oracle grounding further improves the performance substantially, reaching a remarkable whole task success rate of %. Although GPT-4 shows much worse performance than FLAN-T5-XL in step success rate under offline evaluation (Table 2), they exhibit similar success rates in online evaluation. These results further confirm the potential of large models for generalist web agents compared with fine-tuned small models.
@@ 4.3 Further Analysis and Discussion
Error Analysis in Grounding via Image Annotation. Set-of-mark prompting [46] uses a similar method as grounding via image annotation and has been shown effective on object- or scene-centric images [24, 35, 57]. However, this grounding method is suboptimal on webpage screenshot images that are complex and contain rich semantic and spatial relationships.
To analyze the reasons behind the failures, we randomly sample predicted actions and observe the major types of errors as: (1) Wrong action generation; (2) Making up bounding box & label; (3) Failure to link bounding boxes with the correct labels. Illustrative examples are included in Appendix C.
Our analysis reveals that % of the errors originate from wrong action generation wherein the model fail to predict the target element within the action description. This error rate is % higher than images without annotation, which exhibits a % error rate because image annotation often obscures critical content within webpages. Overlaying image annotations on webpages without obscuring important content poses a significant challenge. This is primarily due to the dense layout of webpages, which often feature small elements designed for precise cursor interactions. In contrast, the arrangement of objects in mobile UIs and natural scenes tend to be sparser.
@@
Then % of the errors can be attributed to GPT-4V’s tendency of visual illusion [17], where the model misinterprets and fabricates content over the image. Specifically, the target element described in action generation does not have a bounding box or a label on the bottom-left, where the model is supposed to generate "None". However, the model falsely assumes the presence of a bounding box and makes up a label as the answer. Another % of errors are caused by GPT-4V’s limitation in recognizing the relative position within an image. Specifically, the model is capable of identifying both the target element within the bounding box and the labels. However, it struggles to link the bounding box with its corresponding label.
@@
Powerful Web Action Planner. GPT-4V exhibits promising capabilities, ranging from long-range action planning, webpage content reasoning, and error correction to surpassing the limitations of superficial textual similarity matching inherent in fine-tuned, text-only models. GPT-4V exhibits capability in long-range planning involving a series of interconnected low-level actions. As illustrated in Appendix D, besides generating the immediate next step action, the model also anticipates a sequence of subsequent actions in the future to complete the given task.
Error Correction Awareness. GPT-4V also exhibits the awareness of error correction in the previous actions. In the example in Appendix G, GPT-4V is capable of discovering the mobile phone number is invalid due to wrong format and generate the description about the action to correct this error. This highlights the model’s potential for adaptation in online settings, where actions may not always follow pre-defined, ideal paths as in offline evaluations. This error-correction capability paves the way for adding robustness and reasonable dynamic planning.
@@
Advantages of Knowledge. GPT-4V demonstrates substantial advantages in tasks that requires certain knowledge over fine-tuned models at a smaller scale. As shown in Appendix F, GPT-4V is able to identify the IATA code of the airport in Los Cabos as SJD. In contrast, models are typically weaker at knowledge intensive tasks and also likely to lose knowledge during the fine-tuning process due to catastrophic forgetting.
@@ 5 Related Work
5.1 Large Multimodal Models
@@
The development of LMMs has been characterized by large-scale image-text pretraining to achieve in-context learning capabilities [44, 43, 8, 9, 22, 3, 13, 30, 1, 52]. Further work focuses on improving instruction following [32, 33] capabilities with curated multimodal instruction-following data [59, 11, 26] and alignment with reinforcement learning from human feedback [42]. GPT-4V [31] and Gemini [4] represent significant progress in LMMs. Several studies [2, 31, 49, 54, 46, 45] have highlighted their remarkable multimodal capabilities, emphasizing the advanced and versatile integration of visual and language reasoning abilities. Their performance on a series of benchmarks [21, 15, 19, 37, 28, 56, 53] also showcases remarkable capabilities on vision-and-language understanding and reasoning. Although open-sourced models still exhibit a performance gap with GPT-4V, they have the advantages of controllability and ease of fine-tuning for various applications. For example, in CogAgent [20], LMMs are fine-tuned on HTML and screenshot image pairs to enhance webpage understanding ability and further enhanced with an image encoder for high-resolution image details. Ferret [51] is finetuned to allow visual referring and grounding.
5.2 Web Agent
Driven by the vision of achieving a seamless human-web interaction, considerable initiatives have been invested in web agents that can understand language instruction and complete tasks on the web. Significant efforts have been invested in building benchmarks for web agents with simplified environment simulation [25, 39, 29, 50]. WebArena [58] creates website simulations in the sandbox from four popular categories with functionality and data mimicking their real-world equivalents. Mind2Web [12] instead provides environments in diverse domains, websites, and various tasks on real-world websites by dumping webpages from live websites along with action trajectory annotation. Similar benchmarks have also been proposed for mobile UI automation [5, 23, 41]. To improve web agent performance, WebAgent [18] pretrains a T5 model on HTML data and performs instruction tuning to enable long HTML document summarization and action planning. WebGUM [14] enables web agents to observe both HTML and the rendered screenshot by fine-tuning LMM with a large multimodal corpus for web agents. CogAgent [20] further improves with a high-resolution image encoder to enable the agent to recognize small webpage elements and texts. In a concurrent work [45], GPT-4V exhibits strong performance on mobile UI understanding, which is much simpler than the desktop websites we study.
5.3 Visual Grounding
@@
While LMMs, especially GPT-4V and Gemini, have achieved remarkable vision-language understanding capabilities, they still face challenges in fine-grained visual grounding. Various visual prompting methods have been proposed to augment GPT-4V’s image detail grounding ability. Typically, these methods involve overlaying visual marks onto images, facilitating the model’s reference and grounding to specific regions. Various types of visual marks have been explored, such as red circles [40], mask areas [47], and combinations of circles, arrows, and hand drawings [48]. SoM [46] involves segmenting the image into semantically meaningful regions and overlaying an array of visual marks like numbers, alphabets, masks, or bounding boxes. Fine-tuning vision-language models with image-annotated data has been shown to be effective. Kosmos-2 [34] represents bounding box locations through textual location tokens. BuboGPT [55] extract entities and find corresponding masks for objects in the image. Shikra [7] handles image detail referring and grounding by applying spatial coordinates as text tokens in inputs and outputs, respectively. Ferret [51] represents regions with both discrete coordinates and continuous features along with a spatial-aware visual sampler to handle diverse spatial characteristics across various shapes.
6 Limitations and Potential Societal Impact
Web agents demonstrate a great promise in automating routine web tasks through natural language commands, thereby enhancing web accessibility, especially for people less familiar with information technology. However, there are still potential concerns and limitations regarding our experiment scale and safety concerns for deployment in the real world.
Limited Experiment Scale. We can only run experiments on a subset of Mind2Web due to the limited GPT-4V inference quota. We will seek to expand the scale of the experiments.
Agent Safety. While a generalist web agent holds the potential to automate routine web tasks, enhance user experiences and promote web accessibility, safety concerns related to their real-world deployment are also critical. These concerns span privacy issues, such as access to users’ personal profiles, and sensitive operations, such as financial transactions or application form submissions. During online evaluation, we noticed the possibility for these web agents to generate harmful actions on the web, and we manually validate the safety of all the actions before execution. It is critical for further research to thoroughly assess and mitigate the safety risks associated with web agents, ensuring they are safeguarded against producing and executing harmful actions.
@@ 7 Conclusion
In this work, we developed SeeAct, a generalist web agent that harnesses the power of large multimodal models (LMMs) like GPT-4V to integrate visual understanding and acting on the web. We showed that LMMs present a great promise for generalist web agents, with a success rate of % on live websites given an oracle grounding method. GPT-4V also exhibits impressive capabilities, such as error correction and long-range planning. However, fine-grained visual grounding is still a major challenge. The most effective grounding strategies we explored in this paper still exhibit a -% performance gap compared to oracle grounding. Future work should better leverage the unique properties of the Web, e.g., the known correspondence between HTML and visual elements, for improving grounding and reducing hallucinations from LMMs.
Furthermore, we show a significant discrepancy between online and offline evaluations, emphasizing the importance of online evaluation for an accurate assessment of a model’s capabilities. This discrepancy is largely due to the variability in potential plans for completing the same task, pointing to the dynamic nature of web interactions.
@@ Acknowledgments
Appendix A Offline Experiment Details
@@
The prompt for action generation is shown in Table 1. For grounding via textual choices, image annotation, and element attributes, the prompts are shown in Tabs.˜2, 3 and 4, along with specific task and examples in Figs.˜1, 2, 3, 4 and 5.
@@
| System Role | Imagine that you are imitating humans doing web navigation for a task step by step. At each stage, you can see the webpage like humans by a screenshot and know the previous actions before the current step decided by yourself through recorded history. You need to decide on the first following action to take. You can click an element with the mouse, select an option, or type text with the keyboard. (For your understanding, they are like the click(), select_option() and type() functions in playwright respectively) One next step means one operation within the three. |
| Action Generation | You are asked to complete the following task: {TASK} |
| Previous Actions: {PREVIOUS ACTIONS} | |
@@
| The screenshot below shows the webpage you see. Follow the following guidance to think step by step before outlining the next action step at the current stage: | |
| (Current Webpage Identification) Firstly, think about what the current webpage is. | |
@@
| (Previous Action Analysis) Secondly, combined with the screenshot, analyze each step of the previous action history and their intention one by one. Particularly, pay more attention to the last step, which may be more related to what you should do now as the next step. | |
@@
| (Screenshot Details Analysis) Closely examine the screenshot to check the status of every part of the webpage to understand what you can operate with and what has been set or completed. You should closely examine the screenshot details to see what steps have been completed by previous actions even though you are given the textual previous actions. Because the textual history may not clearly and sufficiently record some effects of previous actions, you should closely evaluate the status of every part of the webpage to understand what you have done. | |
@@
| (Next Action Based on Webpage and Analysis) Then, based on your analysis, in conjunction with human web browsing habits and the logic of web design, decide on the following action. And clearly outline which element in the webpage users will operate with as the first next target element, its detailed location, and the corresponding operation. | |
@@
| To be successful, it is important to follow the following rules: 1. You should only issue a valid action given the current observation. 2. You should only issue one action at a time |
| System Role | Same as Table 1 |
| Action Generation | Slightly modified from Table 1 |
@@
| (Next Action Based on Webpage and Analysis) Then, based on your analysis, in conjunction with human web browsing habits and the logic of web design, decide on the following action. And clearly outline which element in the webpage users will operate with as the first next target element, its detailed location, and the corresponding operation. Please also closely examine the screenshot to adequately describe its position relative to nearby elements and its textual or visual content (if it has). If you find multiple elements similar to your target element, use a more precise description to ensure people can distinguish your target element from them through your answer. | |
@@
| Format Answer | (Final Answer) Finally, conclude your answer using the format below. Ensure your answer is strictly adhering to the format provided below. Please do not leave any explanation in your answers of the final standardized format part, and this final part should be clear and certain. The element, element type, element text, action and value should be in five separate lines. |
| Format: | |
@@
| ELEMENT: Please describe which element you need to operate with. Describe it as detailed as possible, including what it is and where it is. | |
@@
| ELEMENT TYPE: Please specify its type from these options: BUTTON, TEXTBOX, SELECTBOX, or LINK. | |
@@
| ELEMENT TEXT: Please provide the exact text displayed on the element. Do not invent or modify the text; reproduce it as-is from the screenshot. | |
| ACTION: Choose an action from CLICK, TYPE, SELECT. | |
| VALUE: Provide additional input based on ACTION. | |
@@
| System Role | Same as Table 1 |
| Action Generation | Same as Table 1 |
| Referring Description | (Reiteration) First, reiterate your next target element, its detailed location, and the corresponding operation. |
| (Multichoice Question) Below is a multi-choice question, where the choices are elements in the webpage. From the screenshot, find out where and what each one is on the webpage. Then, determine whether one matches your target element. Please examine the choices one by one. Choose the matching one. If multiple options match your answer, choose the most likely one by re-examining the screenshot, the choices, and your further reasoning. | |
| If none of these elements match your target element, please select [None of the other options match the correct element]. A. [CHOICE A] B. [CHOICE B] … | |
| Format Answer | (Final Answer) Finally, conclude your answer using the format below. Ensure your answer is strictly adhering to the format provided below. Please do not leave any explanation in your answers of the final standardized format part, and this final part should be clear and certain. The element choice, action, and value should be in three separate lines. |
| Format: | |
| ELEMENT: The uppercase letter of your choice. | |
| ACTION: Choose an action from CLICK, TYPE, SELECT. | |
| VALUE: Provide additional input based on ACTION. | |
@@
| System Role | Same as Table 1 |
| Action Generation | Slightly modified from Table 1 |
| The screenshot below shows the webpage you see. In the screenshot, some red bounding boxes and white-on-black uppercase letters at the bottom left corner of the bounding boxes have been manually added. You should ignore them for now. Follow the following guidance to think step by step before outlining the next action step at the current stage: | |
| Referring Description | (Reiteration) First, reiterate your next target element, its detailed location, and the corresponding operation. |
| (Verification with the Screenshot) Then, please closely re-examine the screenshot to find whether your target element is marked by a red bounding box and has a white uppercase letter on a black background at the bottom left corner of the bounding box, which is positioned closely next to the bounding box. If yes, use that letter for your final answer. If not, please do not make them up. If it is not marked, please output "NA" as your target element in the following final answer part. | |
| Format Answer | (Final Answer) Finally, conclude your answer using the format below. Ensure your answer is strictly adhering to the format provided below. Please do not leave any explanation in your answers of the final standardized format part, and this final part should be clear and certain. The element choice, action, and value should be in three separate lines. |
| Format: | |
| ELEMENT: The uppercase letter of your choice. | |
| ACTION: Choose an action from CLICK, TYPE, SELECT. | |
| VALUE: Provide additional input based on ACTION. | |
@@
Appendix B Online Experiment Details
We develop an online evaluation tool using Playwright to load webpages and conduct operations generated by web agents. We manually monitor each step of the model and assess whether it finishes the tasks. We explicitly prohibit attempts to log in or perform final submissions to prevent potentially harmful effects.
MindAct. We strictly follow the original settings in MindAct-FLAN-T5 and MindAct-GPT4. The action space only contains Click, Type, and Select.
SeeActOracle We manually implemented the model’s intended actions. The action history was automatically generated by the model, with an added requirement of summarizing the actions into the "Element", "Operation", "Value" format.
SeeActChoice We still adopt the top- and batch them into three option groups as described in offline experiments. We allow PRESS ENTER and TERMINATE for the model to make confirmation, and stop the process.
During our tests, pop-up ads on webpages were manually closed. The MindAct model was not trained on pop-up ads and hence lacks the feature to automatically manage ads, which could result in stalling. Conversely, SeeActcould proactively suggest ad closure through visual analysis and reasoning.
Appendix C Error Examples for Grounding via Image Annotation
In grounding via image annotation method, we observe significant hallucination errors that can be classified into the following categories:
Making up bounding box & label. In our grounding method, if the correct element is absent from the set of candidate elements, the model is anticipated to generate "None" as the answer. However, as depicted in Figure 6 and Figure 7, the model erroneously claims the element is included within a red bounding box and makes up a wrong index label as the answer.
Failure to link bounding boxes with the correct labels. Another challenge arises in accurately linking bounding boxes to their corresponding index labels. This challenge can be attributed to both LMMs’ limitations in understanding relative spatial positions and the complex, dense layout of webpage elements. The model often mistakenly associates the labels of adjacent elements (as illustrated in Figure 8 and Figure 9) or makes up a label entirely (as demonstrated in Figure 10), rather than accurately predicting the intended index label for the targeted element.
Appendix D Strong Capability of Planning
GPT-4V shows remarkable understanding and planning capabilities during our experiments. As depicted in Figure 11, the model is capable of understanding the website and generate a full plans for the given task involving multiple low-level tasks. Specifically, GPT-4V could understand reasonably well about the process and the remaining work of the task by its careful examination of the webpage, as shown in Figure 12.
Appendix E Challenges in Grounding via Textual Choices

### v2

@@ 4.2 Online Evaluation Results
Across all three settings, SeeActChoice with GPT-4V outperforms both GPT-4 and FLAN-T5-XL by a large margin of over 20% whole task success rate. Using oracle grounding further improves the performance substantially, reaching a remarkable whole task success rate of %. Although GPT-4 shows much worse performance than FLAN-T5-XL in step success rate under offline evaluation (Table 2), it outperforms FLAN-T5-XL by 4.4% whole task success rate in the online evaluation. These results further confirm the potential of large models for generalist web agents compared with fine-tuned small models.
4.3 Analysis
@@
Error Analysis in Grounding via Image Annotation. Set-of-mark prompting (Yang et al., 2023a) uses a similar method as grounding via image annotation and has been shown effective on object- or scene-centric images (Lin et al., 2014; Plummer et al., 2015; Zhou et al., 2017). However, this grounding method is suboptimal on webpage screenshot images that are complex and contain rich semantic and spatial relationships. To analyze the reasons behind the failures, we randomly sample action predictions with correct action generation but wrong grounding results. We observes major types of errors as : (1) Making up bounding box & label; (2) Failure to link bounding boxes with the correct labels. Illustrative examples are included in Appendix E.
@@
Our analysis reveals that % of the errors can be attributed to GPT-4V’s tendency of visual illusion (Guan et al., 2023), where the model misinterprets and fabricates content over the image. Specifically, the target element described in action generation does not have a bounding box or a label on the bottom-left, where the model is supposed to generate "NA". However, the model falsely assumes the presence of a bounding box and makes up a label as the answer. Another % of errors are caused by GPT-4V’s limitation in recognizing the relative position within an image. Specifically, the model is capable of identifying the target element within the bounding box. However, it struggles to correctly link the bounding box with its corresponding label.
4.4 Case Study
@@
GPT-4V exhibits promising capabilities, ranging from speculative planning, webpage content reasoning, and error correction to surpassing the limitations of superficial textual similarity matching inherent in fine-tuned, text-only models.
@@
World Knowledge. GPT-4V demonstrates substantial advantages in tasks requiring certain knowledge over fine-tuned models at a smaller scale. As shown in Appendix H, GPT-4V is able to identify the IATA code of the airport in Los Cabos as SJD. In contrast, smaller models are typically weaker at knowledge-intensive tasks and are also likely to lose knowledge during the fine-tuning process due to catastrophic forgetting.
World Model (for Websites). GPT-4V exhibits the potential of a "world model" for websites. As shown in Appendix F, GPT-4V can predict the state transitions on a website (e.g., what would happen if I clicked this button). Based on its awareness of website state transitions, GPT-4V can conduct speculative planning involving a sequence of subsequent actions in the future to complete the given task.
Error Correction Awareness. GPT-4V also exhibits the awareness of error correction in the previous actions. In the example in Appendix I, it realizes that the mobile phone number is invalid due to the wrong format and generates the description of the action to correct this error. This highlights the model’s potential for adaptation in online settings, where actions may not always follow pre-defined, ideal paths as in offline evaluations. This capability paves the way for adding robustness and reasonable dynamic planning.
@@ 5 Related Work
Web Agent. Many works have focused on improving web agents relying on the HTML document (Deng et al., 2023; Gur et al., 2023, 2022, 2023; Kim et al., 2023; Sridhar et al., 2023). However, a raw HTML document is often massive making it infeasible or cost-prohibitively to feed into LLMs directly. MindAct (Deng et al., 2023) instead employs a small language model to rank each HTML element and selectively consider top elements as the context. WebAgent (Gur et al., 2023) proposes an enhanced planning strategy by summarizing the HTML documents and decomposing the instruction into multiple sub-instructions. Another stream considers visual information for web agents (Shaw et al., 2023; Furuta et al., 2023; Hong et al., 2023). Pix2Act (Shaw et al., 2023) leverages Pix2Struct (Lee et al., 2022) to parse screenshot images into simplified HTML to complete GUI-based tasks (Shaw et al., 2023; Liu et al., 2018; Shi et al., 2017; Mazumder & Riva, 2020; Yao et al., 2022). WebGUM (Furuta et al., 2023) and CogAgent (Hong et al., 2023) pre-train an LMM with massive screenshot-HTML data to enhance its decision-making on real-world web navigation like Mind2Web. While all these prior works show promise, generalizing to various web environments remains a challenge for existing models. Thus, SeeAct explores recently released, more powerful LMMs such as GPT-4V and Gemini, to demonstrate their potential as generalist web agents with comprehensive online and offline evaluation and analysis. In a concurrent work (Yan et al., 2023), GPT-4V exhibits strong performance on mobile UI understanding, which is less complex than the desktop websites we study.
@@
Large Multimodal Models. GPT-4V (OpenAI, 2023) and Gemini (Anil et al., 2023) represent significant progress in LMMs. Several studies (Akter et al., 2023; OpenAI, 2023; Yang et al., 2023c; Zhang et al., 2023; Yang et al., 2023a; Yan et al., 2023) have highlighted their remarkable multimodal capabilities, emphasizing the advanced and versatile integration of visual and language reasoning abilities. Their performance on a series of benchmarks (Kazemzadeh et al., 2014; Goyal et al., 2016; Hendrycks et al., 2020; Saikh et al., 2022; Lu et al., 2022; Zhong et al., 2023; Yue et al., 2023) also showcases remarkable capabilities on vision-and-language understanding and reasoning. Although open-sourced models still exhibit a performance gap with GPT-4V, they have the advantages of controllability and ease of fine-tuning for various applications. For example, in CogAgent (Hong et al., 2023), LMMs are fine-tuned on HTML and screenshot image pairs to enhance webpage understanding ability and further enhanced with an image encoder for high-resolution image details. Ferret (You et al., 2023) is finetuned to allow visual referring and grounding.
@@
Visual Grounding. Despite LMMs having achieved remarkable vision-language understanding capabilities, they still face challenges in fine-grained visual grounding. Various visual prompting (Shtedritski et al., 2023; Yang et al., 2023b, c; Yan et al., 2023) methods have been proposed to augment GPT-4V’s image detail grounding ability by overlaying visual marks onto the image. SoM (Yang et al., 2023a) involves segmenting the image into semantically meaningful regions and overlaying an array of visual marks like numbers, alphabets, masks, or bounding boxes. Fine-tuning vision-language models with image-annotated data is effective. Kosmos-2 (Peng et al., 2023) represents bounding box locations through textual location tokens. BuboGPT (Zhao et al., 2023) extract entities and find corresponding masks for objects in the image. Shikra (Chen et al., 2023) handles image detail referring and grounding by applying spatial coordinates as text tokens in inputs and outputs, respectively. Ferret (You et al., 2023) represents regions with both discrete coordinates and continuous features along with a spatial-aware visual sampler to handle diverse spatial characteristics across various shapes.
@@ 6 Conclusion
In this work, we developed SeeAct, a generalist web agent that harnesses the power of large multimodal models (LMMs) like GPT-4V to integrate visual understanding and acting on the web. We showed that LMMs present a great promise for generalist web agents, with a success rate of % on live websites given an oracle grounding method. GPT-4V also exhibits impressive capabilities, such as error correction and speculative planning. However, fine-grained visual grounding is still a major challenge. The most effective grounding strategies we explored in this paper still exhibit a -% performance gap compared to oracle grounding. Future work should better leverage the unique properties of the Web, e.g., the known correspondence between HTML and visual elements, for improving grounding and reducing hallucinations from LMMs. Furthermore, we show a significant discrepancy between online and offline evaluations, emphasizing the importance of online evaluation for an accurate assessment of a model’s capabilities. This discrepancy is largely due to the variability in potential plans for completing the same task, pointing to the dynamic nature of web interactions.
7 Impact Statements
Generalist web agents hold the potential to automate routine web tasks, enhance user experiences, and promote web accessibility, safety concerns related to their real-world deployment are also critical. These concerns span privacy issues, such as access to users’ personal profiles, and sensitive operations, such as financial transactions or application form submissions. During the online evaluation, we noticed the possibility for these web agents to generate harmful actions on the web, and we manually validated the safety of all the actions before execution. It is critical for further research to thoroughly assess and mitigate the safety risks associated with web agents, ensuring they are safeguarded against producing and executing harmful actions. The code will also be released solely for research purposes, with the goal of making the web more accessible via language technologies under an OPEN-RAIL License. We are strongly against any potentially harmful use of the data or technology by any party.
@@ Acknowledgments
Appendix A: Offline Experiments Method Details
2.
Appendix B: Markup Type Ablation Study
3.
LABEL:{appendix:online-exp-details}: Online Experiment Details
4.
Appendix D:Offline Experiment Prompts
5.
Appendix E: Error Examples for Grounding via Image Annotation
6.
Appendix F: Strong Capability of Planning
7.
Appendix G: Challenges in Grounding via Textual Choices
8.
Appendix H: Knowledge and Reasoning Requirements
9.
Appendix I: Path Variation and Awareness of Error Correction
Appendix A Offline Experiments Method Details
FLAN-T5. We fine-tune FLAN-T5 using a left-to-right language modeling objective with the target sequence of ground-truth actions in the Mind2Web training data. The fine-tuned FLAN-T5 then serves as the backbone for inference, enabling action generation in the target format for parsing.
BLIP-2-T5. The BLIP-2 model combines a vision encoder and an LLM with a bridging component for modality connection. We jointly fine-tune the LLM and the bridge module on Mind2Web training data while keeping the vision encoder frozen. For the vision encoder, we leverage the ViT-L/14 pre-trained from CLIP (Radford et al., 2021) with an image resolution of . To ensure a fair comparison with the FLAN-T5-based text-only model, we choose FLAN-T5 as the language model and initialize it with the parameters fine-tuned on Mind2Web.
GPT-3.5 and GPT-4. We also conduct experiments with text-only LLMs, specifically GPT-3.5-turbo-0613 and GPT-4-turbo-1106-preview, using in-context learning in 3-shot settings. We use the same multiple-choice formulation and include three demonstration examples for in-context learning as specified in MindAct.
SeeAct We experiment with GPT-4-vision-preview, Gemini Pro Vision, LLaVA-1.5. Gemini Pro Vision supports only single-turn conversations; therefore, we merge the two turns used in other models for compatibility.
CogAgent We utilize the cogagent-chat-hf checkpoint that hasn’t been fine-tuned on Mind2Web for experiments.
Appendix B Markup Type Ablation Study
The markup types might influence model performance as shown in Yang et al. (2023a); Yan et al. (2023); Koh et al. (2024). We first tested grounding via image annotation through different types of text labels of numerical value, single-digit characters, and two-digit characters, at two different positions to choose a relatively better markup type. The results are shown in Table 5.
| Label | Location | Ele. Acc | Op. F1 | Step SR | |
| Number | Bottom-Left | ||||
| Bottom-Center | |||||
[…]
| Bottom-Center | |||||
[…]
| Bottom-Center | |||||
[…]
Appendix C Online Experiment Details
We develop an online evaluation tool using Playwright to load webpages, acquire textual representation of interactive elements, and perform operations generated by web agents. We manually monitor each step of the model and assess whether it finishes the tasks. Attempts to log in, make final submissions, or perform other potentially harmful actions are prohibited to avoid negative consequences.
MindAct. We adhere strictly to the original settings in MindAct-FLAN-T5 and MindAct-GPT-4, as in offline experiment. For consistency with the MindAct framework and models, we use the scripts from Mind2Web for processing webpages, elements, and generating options, employing the same action space: Click, Type, and Select.
SeeActOracle. In the oracle setting, we manually implement the model’s intended actions. The action history is automatically generated by the model, with an added requirement to summarize actions in the "Element", "Operation", "Value" format in another turn conversation. To avoid overly long screenshots, we use screenshots of current views, and hence allow intentions of scrolling.
SeeActChoice. For a relatively fair comparison, we still adopt the ranker and top- candidate setting, batching them into three option groups as described in offline experiments. We enable PRESS ENTER and TERMINATE for the model to make confirmation or stop the process.
During our tests, pop-up ads on webpages were manually closed. The MindAct model, not trained on handling pop-up ads, lacks the feature to automatically manage them, potentially causing stalls. In contrast, SeeAct models can proactively suggest closing ads through visual analysis and reasoning.
Appendix D Offline Experiment Prompts
@@
The prompt for action generation is shown in Table 6. For grounding via textual choices, image annotation, and element attributes, the prompts are shown in Tables 7, 8 and 9, along with specific tasks and examples in Figures 4, 5, 6, 7 and 8.
| System Role |
@@
Imagine that you are imitating humans doing web navigation for a task step by step. At each stage, you can see the webpage like humans by a screenshot and know the previous actions before the current step decided by yourself through recorded history. You need to decide on the first following action to take. You can click an element with the mouse, select an option, or type text with the keyboard. (For your understanding, they are like the click(), select_option() and type() functions in playwright respectively) One next step means one operation within the three. |
| Action Generation |
You are asked to complete the following task: {TASK} |
Previous Actions: {PREVIOUS ACTIONS} |
@@
The screenshot below shows the webpage you see. Follow the following guidance to think step by step before outlining the next action step at the current stage: |
(Current Webpage Identification) Firstly, think about what the current webpage is. |
@@
(Previous Action Analysis) Secondly, combined with the screenshot, analyze each step of the previous action history and their intention one by one. Particularly, pay more attention to the last step, which may be more related to what you should do now as the next step. |
@@
(Screenshot Details Analysis) Closely examine the screenshot to check the status of every part of the webpage to understand what you can operate with and what has been set or completed. You should closely examine the screenshot details to see what steps have been completed by previous actions even though you are given the textual previous actions. Because the textual history may not clearly and sufficiently record some effects of previous actions, you should closely evaluate the status of every part of the webpage to understand what you have done. |
@@
(Next Action Based on Webpage and Analysis) Then, based on your analysis, in conjunction with human web browsing habits and the logic of web design, decide on the following action. And clearly outline which element in the webpage users will operate with as the first next target element, its detailed location, and the corresponding operation. |
@@
To be successful, it is important to follow the following rules: 1. You should only issue a valid action given the current observation. 2. You should only issue one action at a time. |
| System Role |
Same as Table 6 |
| Action Generation |
Slightly modified from Table 6 |
@@
(Next Action Based on Webpage and Analysis) Then, based on your analysis, in conjunction with human web browsing habits and the logic of web design, decide on the following action. And clearly outline which element in the webpage users will operate with as the first next target element, its detailed location, and the corresponding operation. Please also closely examine the screenshot to adequately describe its position relative to nearby elements and its textual or visual content (if it has). If you find multiple elements similar to your target element, use a more precise description to ensure people can distinguish your target element from them through your answer. |
| Format Answer |
@@
(Final Answer) Finally, conclude your answer using the format below. Ensure your answer is strictly adhering to the format provided below. Please do not leave any explanation in your answers of the final standardized format part, and this final part should be clear and certain. The element, element type, element text, action and value should be in five separate lines. |
Format: |
@@
ELEMENT: Please describe which element you need to operate with. Describe it as detailed as possible, including what it is and where it is. |
@@
ELEMENT TYPE: Please specify its type from these options: BUTTON, TEXTBOX, SELECTBOX, or LINK. |
@@
ELEMENT TEXT: Please provide the exact text displayed on the element. Do not invent or modify the text; reproduce it as-is from the screenshot. |
ACTION: Choose an action from {CLICK, TYPE, SELECT}. |
VALUE: Provide additional input based on ACTION. |
@@
| System Role |
Same as Table 6 |
| Action Generation |
Same as Table 6 |
| Referring Description |
(Reiteration) First, reiterate your next target element, its detailed location, and the corresponding operation. |
(Multichoice Question) Below is a multi-choice question, where the choices are elements in the webpage. From the screenshot, find out where and what each one is on the webpage. Then, determine whether one matches your target element. Please examine the choices one by one. Choose the matching one. If multiple options match your answer, choose the most likely one by re-examining the screenshot, the choices, and your further reasoning. |
If none of these elements match your target element, please select [None of the other options match the correct element]. A. [CHOICE A] B. [CHOICE B] … |
| Format Answer |
(Final Answer) Finally, conclude your answer using the format below. Ensure your answer is strictly adhering to the format provided below. Please do not leave any explanation in your answers of the final standardized format part, and this final part should be clear and certain. The element choice, action, and value should be in three separate lines. |
Format: |
ELEMENT: The uppercase letter of your choice. |
ACTION: Choose an action from {CLICK, TYPE, SELECT}. |
VALUE: Provide additional input based on ACTION. |
@@
| System Role |
Same as Table 6 |
| Action Generation |
Same as Table 6 |
| Referring Description |
(Reiteration) First, reiterate your next target element, its detailed location, and the corresponding operation. |
(Verification with the Screenshot) Then, please closely re-examine the screenshot to find whether your target element is marked by a red bounding box and has a white number on a black background at the bottom left corner of the bounding box, which is positioned closely next to the bounding box. If yes, use that number for your final answer. If not, please do not make them up. If it is not marked, please output "NA" as your target element in the following final answer part. |
| Format Answer |
(Final Answer) Finally, conclude your answer using the format below. Ensure your answer is strictly adhering to the format provided below. Please do not leave any explanation in your answers of the final standardized format part, and this final part should be clear and certain. The element choice, action, and value should be in three separate lines. |
Format: |
ELEMENT: The number of your choice. |
ACTION: Choose an action from {CLICK, TYPE, SELECT}. |
VALUE: Provide additional input based on ACTION. |
@@
Appendix E Error Examples for Grounding via Image Annotation
In the method of grounding via image annotation, we observe significant hallucination errors that can be classified into the following categories:
Making up bounding box & label. In our grounding method, if the correct element is absent from the set of candidate elements, the model is anticipated to generate "NA" as the answer. However, as depicted in Figure 10 and Figure 11, the model erroneously claims the element is included within a red bounding box and makes up a wrong index label as the answer.
Failure to link bounding boxes with the correct labels. Another challenge arises in accurately linking bounding boxes to their corresponding index labels. This challenge can be attributed to both LMMs’ limitations in understanding relative spatial positions and the complex, dense layout of webpage elements. The model often mistakenly associates the labels of adjacent elements (as illustrated in Figure 12 and Figure 13), rather than accurately predicting the intended index label for the targeted element.
Appendix F Strong Capability of Planning
GPT-4V shows remarkable understanding and planning capabilities during our experiments. As depicted in Figure 14, the model is capable of understanding the website and generating a full plan for the given task involving multiple low-level tasks. Specifically, GPT-4V could understand reasonably well about the process and the remaining work of the task by its careful examination of the webpage, as shown in Figure 15.
Appendix G Challenges in Grounding via Textual Choices